# Enhanced Feature Learning via Regularisation: Integrating Neural Networks and Kernel Methods

### v1

@@ A.1.2 Proof of Lemma 4
Proof [Lemma 4] Let us assume now that we only consider functions $f$ with support on the ball with centre 00, radius $R$ and norm $\|\cdot\|^{*}$, which we denote $B(0,R)$. Then we can actually consider the functions $g_{w}$ which define $\mathcal{F}_{\infty}$ to belong to :=$\mathcal{H}^{\prime}:=\{g:\mathbb{R}\to\mathbb{R}\dim g(0)=0,\int_{-R}^{R}(g^{\prime}(t))^{2}{\rm d}t\}$, and it is still a RKHS with the same reproducing kernel. Let $f\in\mathcal{F}_{\infty},f=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}\cdot){\rm d}\mu(w)$. Then, because we have restrained on the ball and we are continuous, $f$ is necessarily in $L_{1}(B(0,R))$ (set of integrable functions) and in $L_{2}(B(0,R))$. It has a Fourier decomposition, with
@@ A.2.1 Proof of Lemma 5
Proof [Lemma 5] Our goal is to transform Equation (5). We begin with the following trick for the $m$ particles setting
@@
Fix $(w_{j})_{j\in[m]}$ and $(\beta_{j})_{j\in[m]}$ in Equation (5), yielding the following minimisation problem on the functions $(g_{j})_{j\in[m]}$
@@
$\min_{c\in\mathbb{R},\alpha^{(1)},\ldots,\alpha^{(m)}\in\mathbb{R}^{d}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},\frac{1}{m}\sum_{j=1}^{m}(K^{(w_{j})}\alpha^{(j)})_{i}+c)+\frac{\lambda}{2}\frac{1}{m}\sum_{j=1}^{m}\frac{(\alpha^{(j)})^{\top}K^{(w_{j})}\alpha^{(j)}}{\beta_{j}}.$
@@
$\min_{\alpha^{(1)},\ldots,\alpha^{(m)}\in\mathbb{R}^{d}}\frac{1}{2}\frac{1}{m}\sum_{j=1}^{m}\frac{(\alpha^{(j)})^{\top}K^{(w_{j})}\alpha^{(j)}}{\beta_{j}},$
@@
$\min_{\tilde{w}_{,}\ldots,\tilde{w}_{m}\in\mathbb{R}^{d},c\in\mathbb{R},\beta\in\mathbb{R}^{m}_{+},\alpha\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},(K\alpha)_{i}+c)+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2}\frac{1}{m}\sum_{j=1}^{m}\|\tilde{w}_{j}\|,$
@@ A.2.2 Proof of Lemma 6
Proof [Lemma 6] The proof follows the same steps as the proof of Lemma 5, systematically replacing any $\frac{1}{m}\sum_{j=1}^{m}$ with the appropriate integral over $\mathcal{S}^{d-1}$ with respect to measure $\mu$. Before the change of variables, the problem is
$\min_{\mu\in\mathcal{P}(\mathcal{S}^{d-1}),c\in\mathbb{R},(\beta_{w})_{w}\in\mathbb{R}_{+}^{\mathcal{S}^{d-1}},\alpha\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},(K\alpha)_{i}+c)+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2}\int_{\mathcal{S}^{d-1}}\beta_{w}{\rm d}\mu(w),$
@@
$\min_{(\beta_{w})_{w}\in\mathbb{R}_{+}^{\mathcal{S}^{d-1}},\nu\in\mathcal{P}(\{\beta_{w}w,w\in\mathcal{S}^{d-1}\}),c\in\mathbb{R},\alpha\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},(K\alpha)_{i}+c)+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2}\int_{\mathbb{R}^{d}}\|\tilde{w}\|{\rm d}\nu(\tilde{w}),$
with $K=\int_{\mathbb{R}^{d}}K^{(\tilde{w})}{\rm d}\nu(\tilde{w})$. We can consider the integral over $\mathbb{R}^{d}$ instead of $\{\beta_{w}w,w\in\mathcal{S}^{d-1}\}$ by extending $\nu$ with $\nu(\mathbb{R}^{d}\setminus\{\beta_{w}w,w\in\mathcal{S}^{d-1}\})=0$. This is equivalent to considering the minimum over $\nu\in\mathcal{P}(\mathbb{R}^{d})$ instead of the minimum over $(\beta_{w})_{w\in\mathcal{S}^{d-1}}\in\mathbb{R}_{+}$ and $\nu\in\mathcal{P}(\{\beta_{w}w,w\in\mathcal{S}^{d-1}\})$.
@@
with $K=\int_{\mathbb{R}^{d}}K^{(\tilde{w})}{\rm d}\nu(\tilde{w})$. Learning an optimal $\nu$ yields an optimal $\mu$ by taking ${\rm d}\mu(w)={\rm d}\nu(\{\tilde{w}\in\mathbb{R}^{d}\mid\tilde{w}/\|\tilde{w}\|=w\})$. For ease of exposition in the main text, we replace $\tilde{w}$ by $w$.
@@ A.3.1 Proof of Lemma 7
Proof [Lemma 7] For a fixed $\alpha$, the optimal $c$ is given by $c=\frac{\mathds{1}^{\top}Y}{n}-\frac{\mathds{1}^{\top}K\alpha}{n}$. Substituting this back into the objective function, we obtain
@@ A.3.2 Proof of Lemma 8
Proof [Lemma 8] First, we compute the derivative of $G=\frac{\lambda}{2}\tilde{Y}^{\top}(\tilde{K}+\lambda nI)^{-1}\tilde{Y}$ with respect to $w_{j}$
@@ A.3.3 Proof of Lemma 9
Proof [Lemma 9] We consider each penalty separately.
@@ A.4.1 Proof of Lemma 11
Proof [Lemma 11] Recall that
@@
We start by considering the expectation over $\varepsilon$ only. Using the definitions, we obtain
$\mathbb{E}_{\varepsilon}$ $\bigg{(}\sup_{f\in\mathcal{F}_{\infty},\Omega(f)\leq D}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(x_{i})\bigg{)}$
$=\mathbb{E}_{\varepsilon}\bigg{(}\sup_{|c|\leq D,\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w)\leq D}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\left(c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}x_{i}){\rm d}\mu(w)\bigg{)}\right)$
$=\mathbb{E}_{\varepsilon}\bigg{(}D\frac{1}{n}\left|\sum_{i=1}^{n}\varepsilon_{i}\right|+\sup_{\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w)\leq D}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\int_{\mathcal{S}^{d-1}}\langle g_{w},k^{(B)}_{w^{\top}x_{i}}\rangle{\rm d}\mu(w)\bigg{)}$
$\leq D\frac{1}{\sqrt{n}}+\mathbb{E}_{\varepsilon}\bigg{(}\sup_{\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w)\leq D}\int_{\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle g_{w},k^{(B)}_{w^{\top}x_{i}}\rangle{\rm d}\mu(w)\bigg{)}$
For the second term of the equation right above, we then have equality to
$=\mathbb{E}_{\varepsilon}\bigg{(}\sup_{\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w)\leq D}\int_{\mathcal{S}^{d-1}}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle g_{w},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}{\rm d}\mu(w)\bigg{)}$
$=\mathbb{E}_{\varepsilon}\bigg{(}\sup_{\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w)\leq D}\sup_{w\in\mathcal{S}^{d-1}}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle g_{w},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}\bigg{)}$
$=\mathbb{E}_{\varepsilon}\bigg{(}\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq D}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle g,k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}\bigg{)}$
$=\mathbb{E}_{\varepsilon}\bigg{(}\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq D}\big{|}\langle g,\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}\bigg{)}=\mathbb{E}_{\varepsilon}\bigg{(}\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq D}\langle g,\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}k^{(B)}_{w^{\top}x_{i}}\rangle\bigg{)}$
$=D\mathbb{E}_{\varepsilon}\bigg{(}\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle g,k^{(B)}_{w^{\top}x_{i}}\rangle\bigg{)}=D\mathbb{E}_{\varepsilon}\bigg{(}\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}g(w^{\top}x_{i})\bigg{)}.$
Taking the expectation over the data set on both sides yields the desired result.
@@ A.4.2 Lemma 18 and its Proof
Proof [Lemma 18] By applying the definitions, we obtain

### v2

@@ A.1.2 Proof of Lemma 4
Proof [Proof of Lemma 4] Let us assume now that we only consider functions $f$ with support on the ball with centre 00, radius $R$ and norm $\|\cdot\|^{*}$, which we denote $B(0,R)$. Then we can actually consider the functions $g$ which define $\mathcal{F}_{\infty}$ to belong to :=$\mathcal{H}^{\prime}:=\{g:\mathbb{R}\to\mathbb{R}\dim g(0)=0,\int_{-R}^{R}(g^{\prime}(t))^{2}{\rm d}t\}$, and it is still a RKHS with the same reproducing kernel. Let $f\in\mathcal{F}_{\infty},f=c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\mu(g,w)$. Then, because we have restrained on the ball and we are continuous, $f$ is necessarily in $L_{1}(B(0,R))$ (set of integrable functions) and in $L_{2}(B(0,R))$. It has a Fourier decomposition, with
@@ A.2.1 Proof of Lemma 5
Proof [Proof of Lemma 5] Our goal is to transform Equation (5). We begin with the following trick for the $m$ particles setting
@@
This allows us to rewrite Equation 5 in the following way
$\min_{c\in\mathbb{R},w_{1},\ldots,w_{m}\in\mathcal{S}^{d-1},g_{1},\ldots,g_{m}\in\mathcal{H},\beta\in\mathbb{R}^{m}_{+}}\frac{1}{n}\sum_{i=1}^{n}\ell\bigg{(}y_{i},c+\frac{1}{m}\sum_{j=1}^{m}g_{j}(w_{j}^{\top}x_{i})\bigg{)}+\lambda\frac{1}{2m}\sum_{j=1}^{m}\left(\frac{\|g_{j}\|_{\mathcal{H}}^{2}}{\beta_{j}}+\beta_{j}\right).$
Now if we fix $(w_{j})_{j\in[m]}$ and $(\beta_{j})_{j\in[m]}$ in the equation above, the minimisation problem is equivalent to
@@
$\min_{c\in\mathbb{R},{\color[rgb]{0,0,0}\alpha^{(1)},\ldots,\alpha^{(m)}\in\mathbb{R}^{{\color[rgb]{0,0,0}n}}}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},\frac{1}{m}\sum_{j=1}^{m}(K^{(w_{j})}\alpha^{(j)})_{i}+c)+\frac{\lambda}{2}\frac{1}{m}\sum_{j=1}^{m}\frac{(\alpha^{(j)})^{\top}K^{(w_{j})}\alpha^{(j)}}{\beta_{j}}.$
@@
$\min_{{\color[rgb]{0,0,0}\alpha^{(1)},\ldots,\alpha^{(m)}\in\mathbb{R}^{n}}}\frac{1}{2}\frac{1}{m}\sum_{j=1}^{m}\frac{(\alpha^{(j)})^{\top}K^{(w_{j})}\alpha^{(j)}}{\beta_{j}},$
@@
$\min_{\tilde{w}_{1},\ldots,\tilde{w}_{m}\in\mathbb{R}^{d},c\in\mathbb{R},\alpha\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},(K\alpha)_{i}+c)+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2}\frac{1}{m}\sum_{j=1}^{m}\|\tilde{w}_{j}\|,$
@@ A.2.2 Proof of Lemma 6
Proof [Proof of Lemma 6] The proof follows the same steps as the proof of Lemma 5, systematically replacing any $\frac{1}{m}\sum_{j=1}^{m}$ with the appropriate integral over ×${\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}$ with respect to measure $\mu$. Before the change of variables, the problem is
$\min_{\mu\in\mathcal{P}({\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}),c\in\mathbb{R},(\beta_{{\color[rgb]{0,0,0}g}})_{{\color[rgb]{0,0,0}g}}\in\mathbb{R}_{+}^{{\color[rgb]{0,0,0}\mathcal{H}}},\alpha\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},(K\alpha)_{i}+c)+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2}\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}{\color[rgb]{0,0,0}\beta_{g}}{\rm d}\mu({\color[rgb]{0,0,0}g,w}),$
@@
$\min_{(\beta_{{\color[rgb]{0,0,0}g}})_{{\color[rgb]{0,0,0}g}}\in\mathbb{R}_{+}^{{\color[rgb]{0,0,0}\mathcal{H}}},\nu\in\mathcal{P}(\{\beta_{{\color[rgb]{0,0,0}g}}w,{\color[rgb]{0,0,0}g\in\mathcal{H}},w\in\mathcal{S}^{d-1}\}),c\in\mathbb{R},\alpha\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},(K\alpha)_{i}+c)+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2}\int_{\mathbb{R}^{d}}\|\tilde{w}\|{\rm d}\nu(\tilde{w}),$
with $K=\int_{\mathbb{R}^{d}}K^{(\tilde{w})}{\rm d}\nu(\tilde{w})$. We can consider the integral over $\mathbb{R}^{d}$ instead of $\{\beta_{{\color[rgb]{0,0,0}g}}w,{\color[rgb]{0,0,0}g\in\mathcal{H}},w\in\mathcal{S}^{d-1}\}$ by extending $\nu$ with $\nu(\mathbb{R}^{d}\setminus\{\beta_{{\color[rgb]{0,0,0}g}}w,{\color[rgb]{0,0,0}g\in\mathcal{H}},w\in\mathcal{S}^{d-1}\})=0$. This is equivalent to considering the minimum over $\nu\in\mathcal{P}(\mathbb{R}^{d})$ instead of the minimum over $(\beta_{{\color[rgb]{0,0,0}g}})_{{\color[rgb]{0,0,0}g\in\mathcal{H}}}\in\mathbb{R}_{+}^{{\color[rgb]{0,0,0}\mathcal{H}}}$ and $\nu\in\mathcal{P}(\{\beta_{{\color[rgb]{0,0,0}g}}w,{\color[rgb]{0,0,0}g\in\mathcal{H}},w\in\mathcal{S}^{d-1}\})$, as we discuss now.
@@
with $K=\int_{\mathbb{R}^{d}}K^{(\tilde{w})}{\rm d}\nu(\tilde{w})$. For ease of exposition in the main text, we replace $\tilde{w}$ by $w$.
@@ A.3.1 Proof of Lemma 7
Proof [Proof of Lemma 7] For a fixed $\alpha$, the optimal $c$ is given by $c=\frac{\mathds{1}^{\top}Y}{n}-\frac{\mathds{1}^{\top}K\alpha}{n}$. Substituting this back into the objective function, we obtain
@@ A.3.2 Proof of Lemma 8
Proof [Proof of Lemma 8] First, we compute the derivative of $G=\frac{\lambda}{2}\tilde{Y}^{\top}(\tilde{K}+\lambda nI)^{-1}\tilde{Y}$ with respect to $w_{j}$
@@ A.3.3 Proof of Lemma 9
Proof [Proof of Lemma 9] We consider each penalty separately.
@@ A.4.1 Proof of Lemma 11
Proof [Proof of Lemma 11] Recall that
@@
We start by considering the quantity without any expectation. Using the definitions, we obtain
$\sup_{f\in\mathcal{F}_{\infty},\Omega(f)\leq D}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(x_{i})$ $=\sup_{|c|\leq D,{\color[rgb]{0,0,0}\ \mu\ \text{s.t.}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)}\leq D}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\left(c+{\color[rgb]{0,0,0}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}x_{i}){\rm d}\mu(g,w)}\right)$
$=D\frac{1}{n}\left|\sum_{i=1}^{n}\varepsilon_{i}\right|+\sup_{\ \mu\ \text{s.t.}\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}{\color[rgb]{0,0,0}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)}\leq D}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}\langle{\color[rgb]{0,0,0}g},k^{(B)}_{w^{\top}x_{i}}\rangle{\rm d}{\color[rgb]{0,0,0}\mu(g,w)}$
where the last equality is obtained by splitting the sup and taking $c=D{\rm sign}\left(\sum_{i=1}^{n}\varepsilon_{i}\right)$, which is the explicit value attaining the supremum. For the first term, we can simply bound $\mathbb{E}_{\varepsilon}\bigg{(}D\frac{1}{n}\left|\sum_{i=1}^{n}\varepsilon_{i}\right|\bigg{)}\leq D\frac{1}{\sqrt{n}}$ using the gaussianity of $\varepsilon$. For the second term of the equation right above, we then have equality to
$\sup_{\ \mu\ \text{s.t.}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)\leq D}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle g,k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}{\rm d}\mu(g,w)$
$\leq\sup_{\ \mu\ \text{s.t.}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)\leq D}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\mathds{1}_{g\neq 0}\|g\|_{\mathcal{H}}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle\frac{g}{\|g\|_{\mathcal{H}}},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}{\rm d}\mu(g,w)$
$\leq\sup_{\ \mu\ \text{s.t.}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)\leq D}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\mathds{1}_{g\neq 0}\|g\|_{\mathcal{H}}\sup_{w\in\mathcal{S}^{d-1},\|\tilde{g}\|_{\mathcal{H}}=1}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle\tilde{g},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}{\rm d}\mu(g,w)$
$\leq\sup_{\ \mu\ \text{s.t.}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)\leq D}\sup_{w\in\mathcal{S}^{d-1},\|\tilde{g}\|_{\mathcal{H}}=1}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle\tilde{g},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\mathds{1}_{g\neq 0}\|g\|_{\mathcal{H}}{\rm d}\mu(g,w)$
$\leq D\sup_{w\in\mathcal{S}^{d-1},\|\tilde{g}\|_{\mathcal{H}}=1}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle\tilde{g},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}\leq D\sup_{w\in\mathcal{S}^{d-1},\|\tilde{g}\|_{\mathcal{H}}\leq 1}\big{|}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle\tilde{g},k^{(B)}_{w^{\top}x_{i}}\rangle\big{|}$
$\leq D\sup_{w\in\mathcal{S}^{d-1},\|\tilde{g}\|_{\mathcal{H}}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\langle\tilde{g},k^{(B)}_{w^{\top}x_{i}}\rangle=D\sup_{f=\tilde{g}(w^{\top}\cdot),w\in\mathcal{S}^{d-1},\|\tilde{g}\|_{\mathcal{H}}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(x_{i}).$
Taking the expectation over the data set and $\varepsilon$ on both sides and renaming $\tilde{g}$ as $g$ for ease of exposition yields the desired result.
@@ A.4.2 Lemma 18 and its Proof
Proof [Proof of Lemma 18] By applying the definitions, we obtain